In the laboratory you dissolve 20.9 g of copper(II) iodide in a volumetric flask and add water to a total volume of 500 mL.


What is the molarity of the solution?  _____M.


What is the concentration of the copper(II) cation?  ______M.


What is the concentration of the iodide anion?  ______M.


1
Expert's answer
2021-10-21T02:09:21-0400

"m=20.9 \;g \\
V=500 \; mL =0.5 \;L\\
M(CuI_2) = 317.35 \;g/mol \\
n = \frac{20.9}{317.35} = 0.0658 \;mol \\
Molality = \frac{n}{V} = \frac{0.0658}{0.5} = 0.1316 \;M \\
[CuI_2] = 0.1316 \;M \\
[Cu^{2+}] = [CuI_2] = 0.1316 \;M \\
[I^{-1}] = 2[Cu^{2+}] = 2 \times 0.1316 = 0.2632 \;M"